# arXiv:2602.23381v1[math.GN] 03 Feb 2026

### v1

@@ Abstract
We develop a framework for shallow and deep neural networks whose inputs range over a general topological space. The model is built from a prescribed family of continuous feature maps and a fixed scalar activation function, and it reduces to multilayer feedforward networks in the Euclidean case. We focus on the universal approximation property and establish general conditions under which such networks are dense in spaces of continuous vector-valued functions on arbitrary and locally convex topological spaces. In the absence of width constraints, we obtain universality results that extend classical approximation theorems to non-Euclidean settings. A central focus of the paper is the deep narrow framework, in which the width of each hidden layer is uniformly bounded while the depth is allowed to grow. We identify conditions under which such width-constrained deep networks retain universal approximation power. As a concrete example, we employ Ostrand’s extension of the Kolmogorov superposition theorem to derive an explicit universality result for products of compact metric spaces, with width bounds expressed in terms of topological dimension.
@@ 1 Introduction
Neural networks are a central tool in contemporary machine learning and artificial intelligence. Among the many architectures in use, multilayer feedforward neural networks remain a basic and influential model, both in practice and in theory. A broad overview of the approximation-theoretic properties of feedforward neural networks can be found, for example, in [28, 11], while modern deep learning practice is surveyed in [7].
A multilayer feedforward neural network consists of an input layer, an output layer, and one or more hidden layers arranged sequentially. Information propagates forward through the network: each neuron forms an affine combination of the outputs of the previous layer and then applies a fixed nonlinear activation function. This repeated affine–nonlinear processing enables the network to represent highly complex nonlinear mappings.
@@
Note that the inner product $\mathbf{w}\cdot\mathbf{x}$ appearing in classical neural networks is a linear continuous functional on $\mathbb{R}^{d}$. Conversely, by the Riesz representation theorem, every linear continuous functional on $\mathbb{R}^{d}$ can be written in this form (see [30, Theorem 13.32]). Let $\mathcal{L}(\mathbb{R}^{d})$ denote the space of all linear continuous functionals on $\mathbb{R}^{d}$. From this viewpoint, a single hidden layer network applies a collection of functions from $\mathcal{L}(\mathbb{R}^{d})$ to the input, adds scalar biases, and then applies a nonlinearity. In particular, the classical universal approximation property can be rephrased as the density of the function class
@@
We now define shallow topological feedforward neural networks in a form that directly mirrors the standard affine–activation–affine structure of classical single hidden layer networks.
@@
$H(x)=A\,\sigma\!\bigl(T(x)\bigr)-b,$
@@
  • $A\in\mathbb{R}^{m\times r}$ and $b\in\mathbb{R}^{m}$ are the output-layer weights and bias;
@@
where both $T_{0}$ and $T_{1}$ are affine maps.
@@
$H=T_{l+1}\circ\sigma\circ T_{l}\circ\sigma\circ\cdots\circ\sigma\circ T_{1}\circ\sigma\circ T_{0},$
@@
  • for $j=1,\dots,l$, $T_{j}:\mathbb{R}^{k_{j-1}}\to\mathbb{R}^{k_{j}}$ is an affine map
@@
  • $T_{l+1}:\mathbb{R}^{k_{l}}\to\mathbb{R}^{m}$ is an affine output map
    $T_{l+1}(y)=A_{l+1}y-b_{l+1},\qquad A_{l+1}\in\mathbb{R}^{m\times k_{l}},\;b_{l+1}\in\mathbb{R}^{m};$
@@
This definition mirrors exactly the standard representation of deep feedforward neural networks as compositions of affine maps and componentwise nonlinearities. In particular, when $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, deep TFNNs reduce precisely to classical deep neural networks of the form
$\Phi(x)=T_{l+1}\circ\sigma\circ T_{l}\circ\cdots\circ\sigma\circ T_{0}(x),$
with each $T_{j}$ affine.
@@
Aim of the paper. The aim of this paper is to develop a universality theory for TFNNs with vector-valued outputs on general topological input spaces. More precisely, we seek conditions on the feature family $\mathcal{A}(X)\subset C(X)$ and on the activation function $\sigma$ under which, for every compact set $K\subset X$, deep TFNNs are dense in the space $C(K;\mathbb{R}^{m})$ of continuous $\mathbb{R}^{m}$-valued functions on $K$.
A central focus of the paper is the deep narrow setting, in which the width of each hidden layer is uniformly bounded while the depth is allowed to be arbitrary. Our goal is to identify structural assumptions ensuring that such width-constrained deep TFNNs retain universal approximation power. In this way, we extend the deep narrow universality property known for Euclidean input spaces to a broad class of non-Euclidean spaces.
Related literature. The approximation theory of neural networks is centered around the universal approximation property (UAP), also referred to as density. In the classical Euclidean setting, this property asserts that, for suitable activation functions $\sigma$, neural networks with a single hidden layer are dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$. The most general characterization of this phenomenon was obtained by Leshno, Lin, Pinkus, and Schocken [23], who proved that, under the assumption of continuity, an activation function $\sigma$ has the UAP if and only if it is not a polynomial. This result demonstrates the expressive power of shallow feedforward networks for a broad class of activation functions. It is worth noting that the theorem in [23] applies to a wider class of activation functions, including some that are discontinuous on sets of Lebesgue measure zero. Detailed proofs and further discussion can be found in [27, 28].
@@
A different perspective on neural network approximation over non-Euclidean spaces relies on abstract transfer principles, through which approximation properties established on Euclidean spaces extend to more general domains. For example, [20] investigated which modifications of the input and output layers preserve the universal approximation property. The authors showed that, under suitable conditions on topological spaces $X$ and $Y$ and continuous maps $\phi:X\to\mathbb{R}^{n}$ and $\rho:\mathbb{R}^{m}\to Y$, density of a class $\mathcal{F}\subset C(\mathbb{R}^{n},\mathbb{R}^{m})$ implies density of the composed class $\{\rho\circ f\circ\phi:f\in\mathcal{F}\}$ in $C(X,Y)$ with respect to uniform convergence on compact sets.
@@
It is worth noting that, in the shallow setting, universality for scalar-valued networks on Banach spaces of continuous functions [2] and on general topological spaces [14], as well as for networks with values in finite-dimensional algebras [34], has been established for activation functions belonging to the Tauber–Wiener class. Moreover, for topological spaces with the Hahn–Banach extension property, scalar-valued universality can be characterized by non-polynomiality of the activation function on a sufficiently small interval [13].
The present work differs from much of the above literature in two essential aspects. First, in addition to treating vector-valued shallow (and deep) networks without width constraints, we investigate deep networks with a uniform width bound and study whether universality can be retained in this setting. Second, we consider approximation of vector-valued functions defined on arbitrary topological input spaces. The only restriction of our main result is that the output space must be a finite-dimensional Euclidean space. The techniques developed here do not yet extend to neural networks with outputs in infinite-dimensional spaces. However, we hope that the results of this paper will stimulate further exploration of deep neural networks with outputs taking values in abstract infinite-dimensional spaces.
@@ 2 Universal approximation on topological spaces
Thus, in what follows, when we say that a family of functions $\mathcal{F}$ acting from $X$ into $\mathbb{R}^{m}$ is dense in $C(X;\mathbb{R}^{m})$, we always mean density with respect to the above topology of uniform convergence on compact sets. That is, $\mathcal{F}$ is dense in $C(X;\mathbb{R}^{m})$ if, for every compact set $K\subset X$, every $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists $f\in\mathcal{F}$ such that
@@
Our objective is to relate approximation on the abstract input space $X$ to approximation on Euclidean spaces. This requires, on the one hand, suitable richness assumptions on the feature family $\mathcal{A}(X)\subset C(X)$, and, on the other hand, approximation properties of the activation function $\sigma$. In this context we distinguish between two cases. The first case, which covers shallow and deep TFNNs without width constraints, can be treated using a structural density condition on the feature family, called the $D$-property. The second case, corresponding to deep but narrow networks, requires additional ideas that reduce the approximation problem to a finite-dimensional setting and exploit depth to compensate for width constraints.
2.1 The $D$-property and universality without width constraints
@@
Equivalently, $\mathcal{A}(X)$ has the $D$-property if and only if for every compact set $K\subset X$, every function $g\in C(K)$, and every $\varepsilon>0$, there exist an integer $m\in\mathbb{N}$, feature maps $f_{1},\dots,f_{m}\in\mathcal{A}(X)$, and univariate continuous functions $u_{1},\dots,u_{m}\in C(\mathbb{R})$ such that
@@ 2.2 A univariate approximation assumption
To realize compositions of the form $u\circ f$ by neural networks, we require that the activation function be able to approximate arbitrary continuous functions of a single real variable. The following abstract assumption captures this requirement.
@@
2.3 Universality of TFNNs without width constraints
@@
Let $X$ be a topological space and let $\mathcal{A}(X)\subset C(X)$ be a feature family with the $D$-property. Assume that the activation function $\sigma:\mathbb{R}\to\mathbb{R}$ satisfies the univariate universal approximation property of Definition 2.2. Then for every compact set $K\subset X$, every function $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists a TFNN $H\in\mathcal{N}_{1}^{(m)}(\sigma)$ such that
@@
Consequently, for every $l\geq 1$, the deep TFNN class $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense in $C(X;\mathbb{R}^{m})$.
@@
Fix a compact set $K\subset X$, a vector-valued function
$g=(g_{1},\dots,g_{m})\in C(K;\mathbb{R}^{m}),$
and $\varepsilon>0$. We begin by constructing, for $l=1$, a TFNN $H=(H_{1},\dots,H_{m})$ satisfying
@@
Fix $k\in\{1,\dots,m\}$. The function $g_{k}$ is continuous on the compact set $K$. By the $D$-property of $\mathcal{A}(X)$, the linear span of compositions
@@
is dense in $C(X)$ in the topology of uniform convergence on compact sets. Applying this density to the restriction on $K$, we can approximate $g_{k}$ uniformly on $K$ by a finite sum of such compositions. Concretely, there exist an integer $M_{k}\in\mathbb{N}$, feature maps $f_{k,1},\dots,f_{k,M_{k}}\in\mathcal{A}(X)$, and continuous functions $u_{k,1},\dots,u_{k,M_{k}}\in C(\mathbb{R})$ such that
@@
For the deep classes, note that every shallow network may be viewed as a deep network by inserting identity (or otherwise trivial) affine layers. Formally, by Definition 1.3, we have
$\mathcal{N}_{1}^{(m)}(\sigma)\subset\mathcal{N}_{l}^{(m)}(\sigma),\qquad l\geq 1.$
If a set $B$ of a topological vector space is dense, then any superset of $B$ is also dense. Hence each $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense as well. This completes the proof. ∎

### v2

# Universality of shallow and deep neural networks on non-Euclidean spaces

@@ Abstract
We study shallow and deep neural networks whose inputs range over a general topological space. The model is built from a prescribed family of continuous feature maps and reduces to multilayer feedforward networks in the Euclidean case. We focus on the universal approximation property and establish general conditions under which such networks are dense in spaces of continuous vector-valued functions on arbitrary topological spaces and, in particular, locally convex spaces. Universality results obtained in the arbitrary-width case extend classical approximation theorems to non-Euclidean spaces. We also consider the deep narrow setting, in which the width of each hidden layer is uniformly bounded while the depth is allowed to grow. We identify conditions under which such networks retain the universal approximation property. As a concrete example, we employ Ostrand’s extension of the Kolmogorov superposition theorem to derive an explicit universality result for products of compact metric spaces, with width bounds expressed in terms of topological dimension.
@@ 1 Introduction
Neural networks play a central role in modern machine learning and artificial intelligence. Among the many architectures in use, multilayer feedforward neural networks remain a basic and influential model in both theory and practice. A broad overview of the approximation-theoretic properties of feedforward neural networks can be found, for example, in [28, 11], while modern deep learning practice is surveyed in [7].
A multilayer feedforward neural network consists of an input layer, an output layer, and one or more hidden layers arranged sequentially. Information propagates forward through the network: each neuron forms an affine combination of the outputs of the previous layer and then applies a fixed nonlinear activation function. This repeated affine–nonlinear processing enables the network to represent complex nonlinear mappings.
@@
Note that the inner product $\mathbf{w}\cdot\mathbf{x}$ appearing in classical neural networks is a linear continuous functional on $\mathbb{R}^{d}$. Conversely, by the Riesz representation theorem, every linear continuous functional on $\mathbb{R}^{d}$ can be written in this form (see [30, Theorem 13.32]). Let $\mathcal{L}(\mathbb{R}^{d})$ denote the space of all linear continuous functionals on $\mathbb{R}^{d}$. Under this formulation, a single hidden layer network applies a collection of functions from $\mathcal{L}(\mathbb{R}^{d})$ to the input, adds scalar biases, and then applies a nonlinearity. In particular, the classical universal approximation property can be rephrased as the density of the function class
@@
We now define shallow topological feedforward neural networks in a form that corresponds to the standard affine–activation–affine structure of classical single hidden layer networks.
@@
$H(x)=A\,\sigma\!\bigl(T(x)\bigr),$
@@
  • $A\in\mathbb{R}^{m\times r}$ is the linear output-layer weight matrix;
@@
where $T_{0}$ is affine and $T_{1}$ is linear.
In the case $m=1$, the network in Definition 1.2 can be written in the classical scalar form
$H(x)=\sum_{i=1}^{r}c_{i}\,\sigma\!\bigl(w_{i}f_{i}(x)-\theta_{i}\bigr),$
where $c_{i}\in\mathbb{R}$ are the output weights.
More generally, for $m\geq 1$ the vector-valued network can be written componentwise as
$H_{j}(x)=\sum_{i=1}^{r}c_{ji}\,\sigma\!\bigl(w_{i}f_{i}(x)-\theta_{i}\bigr),\qquad j=1,\dots,m,$
where $A=(c_{ji})\in\mathbb{R}^{m\times r}$ is the output weight matrix.
@@
$H=T_{l}\circ\sigma\circ T_{l-1}\circ\sigma\circ\cdots\circ\sigma\circ T_{1}\circ\sigma\circ T_{0},$
@@
  • for $j=1,\dots,l-1$, $T_{j}:\mathbb{R}^{k_{j-1}}\to\mathbb{R}^{k_{j}}$ is an affine map
@@
  • $T_{l}:\mathbb{R}^{k_{l-1}}\to\mathbb{R}^{m}$ is a linear output map
    $T_{l}(y)=A_{l}y,\qquad A_{l}\in\mathbb{R}^{m\times k_{l-1}};$
@@
This definition agrees with the standard representation of deep feedforward neural networks as compositions of affine maps in the hidden layers, componentwise nonlinearities, and a linear output map. In particular, when $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, deep TFNNs reduce precisely to classical deep neural networks of the form
$\Phi(x)=T_{l}\circ\sigma\circ T_{l-1}\circ\cdots\circ\sigma\circ T_{0}(x),$
with each $T_{j}$ affine for $j=0,\dots,l-1$ and $T_{l}$ linear.
@@
Aim of the paper. The aim of this paper is to develop a universality theory for TFNNs with vector-valued outputs whose inputs lie in an arbitrary topological space $X$. More precisely, we seek conditions on the feature family $\mathcal{A}(X)\subset C(X)$ and on the activation function $\sigma$ under which, for every compact set $K\subset X$, TFNNs are dense in the space $C(K;\mathbb{R}^{m})$ of continuous $\mathbb{R}^{m}$-valued functions on $K$.
We also focus on the deep narrow setting, in which the width of each hidden layer is uniformly bounded while the depth is allowed to be arbitrary. Our goal is to identify structural assumptions ensuring that such bounded-width deep TFNNs retain the universal approximation property. In this way, we extend the deep narrow universality property known for Euclidean input spaces to a broad class of non-Euclidean spaces.
Related literature. The approximation theory of neural networks is centered around the universal approximation property (UAP), also referred to as density. In the classical Euclidean setting, this property asserts that, for suitable activation functions $\sigma$, neural networks with a single hidden layer are dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$. The most general characterization of this phenomenon was obtained by Leshno, Lin, Pinkus, and Schocken [23], who proved that, under the assumption of continuity, an activation function $\sigma$ has the UAP if and only if it is not a polynomial. This result demonstrates the expressive power of shallow feedforward networks for a broad class of activation functions. Note that the theorem in [23] applies to a wider class of activation functions, including some that are discontinuous on sets of Lebesgue measure zero. Detailed proofs and further discussion can be found in [27, 28].
@@
An alternative approach to neural network approximation over non-Euclidean spaces relies on certain transfer principles, through which approximation properties established on Euclidean spaces extend to more general domains. For example, [20] investigated which modifications of the input and output layers preserve the universal approximation property. The authors showed that, under suitable conditions on topological spaces $X$ and $Y$ and continuous maps $\phi:X\to\mathbb{R}^{n}$ and $\rho:\mathbb{R}^{m}\to Y$, density of a class $\mathcal{F}\subset C(\mathbb{R}^{n},\mathbb{R}^{m})$ implies density of the composed class $\{\rho\circ f\circ\phi:f\in\mathcal{F}\}$ in $C(X,Y)$ with respect to uniform convergence on compact sets.
@@
It should be remarked that, in the shallow setting, universality for scalar-valued networks on Banach spaces of continuous functions [2] and on general topological spaces [14], as well as for networks with values in finite-dimensional algebras [34], has been established for activation functions belonging to the Tauber–Wiener class. Moreover, for topological spaces with the Hahn–Banach extension property, scalar-valued universality can be characterized by non-polynomiality of the activation function on a sufficiently small interval [13].
The present work differs from much of the above literature in two essential aspects. First, in addition to treating vector-valued shallow networks without width constraints, we investigate deep networks with a uniform width bound and study whether universality can be retained in this setting. Second, we consider approximation of vector-valued functions defined on arbitrary topological input spaces. The only restriction of our main result is that the output space must be a finite-dimensional Euclidean space. The techniques developed here do not yet extend to neural networks with outputs in infinite-dimensional spaces. However, we hope that the results of this paper will stimulate further exploration of shallow and deep neural networks with outputs taking values in abstract infinite-dimensional spaces.
@@ 2 Universal approximation on topological spaces
Thus, in what follows, when we say that a family of functions $\mathcal{F}$ acting from $X$ into $\mathbb{R}^{m}$ is dense in $C(X;\mathbb{R}^{m})$, we always mean density with respect to the above topology of uniform convergence on compact sets. That is, $\mathcal{F}$ is dense in $C(X;\mathbb{R}^{m})$ if, for every $g\in C(X;\mathbb{R}^{m})$, every compact set $K\subset X$, and every $\varepsilon>0$, there exists $f\in\mathcal{F}$ such that
@@
Our objective is to relate approximation on the abstract input space $X$ to approximation on Euclidean spaces. This requires, on the one hand, suitable richness assumptions on the feature family $\mathcal{A}(X)\subset C(X)$, and, on the other hand, approximation properties of the activation function $\sigma$. In this context we distinguish between two cases. The first case, which covers shallow TFNNs without width constraints, can be treated using a structural density condition on the feature family, called the $D$-property. The second case, corresponding to deep but narrow networks, requires additional ideas that reduce the approximation problem to a finite-dimensional setting and exploit depth to compensate for width constraints.
2.1 The $D$-property
@@
Equivalently, $\mathcal{A}(X)$ has the $D$-property if and only if for every function $g\in C(X)$, every compact set $K\subset X$, and every $\varepsilon>0$, there exist an integer $m\in\mathbb{N}$, feature maps $f_{1},\dots,f_{m}\in\mathcal{A}(X)$, and univariate continuous functions $u_{1},\dots,u_{m}\in C(\mathbb{R})$ such that
@@ 2.2 A univariate approximation assumption
To approximate compositions of the form $u\circ f$ by neural networks, we require that the activation function be able to approximate arbitrary continuous functions of a single real variable. The following assumption captures this requirement.
@@
2.3 Universality of TFNNs in the arbitrary-width case
@@
For a topological space $X$, a feature family $\mathcal{A}(X)\subset C(X)$, and an activation function $\sigma:\mathbb{R}\to\mathbb{R}$, we denote by $\mathcal{N}_{1}^{(m)}(\sigma)$ the class of all TFNNs with one hidden layer and $m$ outputs of the form described in Definition 1.2. More generally, for each $l\geq 1$, we denote by $\mathcal{N}_{l}^{(m)}(\sigma)$ the class of all deep TFNNs of depth $l$ with $m$ outputs as in Definition 1.3.
@@
Let $X$ be a topological space and let $\mathcal{A}(X)\subset C(X)$ be a feature family with the $D$-property. Assume that the activation function $\sigma:\mathbb{R}\to\mathbb{R}$ satisfies the univariate universal approximation property. Then the TFNN class $\mathcal{N}_{1}^{(m)}(\sigma)$ is dense in $C(X;\mathbb{R}^{m})$ with respect to the topology of uniform convergence on compact sets. That is, for every function $g\in C(X;\mathbb{R}^{m})$, every compact set $K\subset X$, and every $\varepsilon>0$, there exists a TFNN $H\in\mathcal{N}_{1}^{(m)}(\sigma)$ such that
@@
Fix a vector-valued function
$g=(g_{1},\dots,g_{m})\in C(X;\mathbb{R}^{m}),$
a compact set $K\subset X$ and $\varepsilon>0$. We construct a TFNN $H=(H_{1},\dots,H_{m})$ satisfying
@@
Fix $k\in\{1,\dots,m\}$ and consider the scalar function $g_{k}$. By the $D$-property of $\mathcal{A}(X)$, the linear span of compositions
@@
is dense in $C(X)$ in the topology of uniform convergence on compact sets. Hence, we can approximate $g_{k}$ uniformly on $K$ by a finite sum of such compositions. Concretely, there exist an integer $M_{k}\in\mathbb{N}$, feature maps $f_{k,1},\dots,f_{k,M_{k}}\in\mathcal{A}(X)$, and continuous functions $u_{k,1},\dots,u_{k,M_{k}}\in C(\mathbb{R})$ such that
@@
This completes the proof. ∎
In Theorem 2.1, the target function is taken from $C(X;\mathbb{R}^{m})$ rather than $C(K;\mathbb{R}^{m})$ for a fixed compact set $K\subset X$. This is due to the formulation of the $D$-property as a density statement in $C(X)$ with respect to the topology of uniform convergence on compact sets.
In general topological spaces, a function defined on a compact subset $K$ need not admit a continuous extension to the whole space $X$. Therefore, working with functions in $C(X;\mathbb{R}^{m})$ avoids the need for extension arguments. However, in spaces where continuous extension from closed sets is available, such as normal topological spaces, one may equivalently formulate the approximation property for functions in $C(K;\mathbb{R}^{m})$.
Corollary 2.1.
Let $X$ be a normal topological space and let $\mathcal{A}(X)\subset C(X)$ satisfy the $D$-property. Assume that $\sigma$ satisfies the univariate universal approximation property. Then for every compact set $K\subset X$, every function $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists $H\in\mathcal{N}_{1}^{(m)}(\sigma)$ such that
$\|g-H\|_{K}<\varepsilon.$
Proof.
Let $K\subset X$ be compact, let $g=(g_{1},\dots,g_{m})\in C(K;\mathbb{R}^{m})$, and let $\varepsilon>0$. Since $X$ is normal and $K$ is closed, the Tietze extension theorem implies that for each $j=1,\dots,m$ there exists $G_{j}\in C(X)$ such that $G_{j}|_{K}=g_{j}$. Define $G=(G_{1},\dots,G_{m})\in C(X;\mathbb{R}^{m})$; then $G|_{K}=g$.
Since $\mathcal{A}(X)$ has the $D$-property and $\sigma$ satisfies the univariate universal approximation property, Theorem 2.1 yields a TFNN $H\in\mathcal{N}_{1}^{(m)}(\sigma)$ such that
$\|G-H\|_{K}<\varepsilon.$
Because $G=g$ on $K$, it follows that $\|g-H\|_{K}=\|G-H\|_{K}<\varepsilon$. ∎